OpenStudy (anonymous):

1. Find the difference quotient for f(x) = 4x^2 - 2x 2. Let f(x)= x^2 - 25 and g(x) = x + 5. Find and simplify a. (f - g)(x) b. (f/g)(x)

OpenStudy (anonymous):

f(x+h)-f(x) ----------- h

OpenStudy (anonymous):

@ odehye can u start it

OpenStudy (anonymous):

Thanks Mark

OpenStudy (anonymous):

I'm starting

OpenStudy (anonymous):

f(x) = 4x^2 - 2x f(x+h) = 4(x+h)^2 - 2(x+h) f(x+h) = 4(x^2+2xh+h^2) - 2(x+h) f(x+h) = 4x^2+8xh+4h^2 - 2x-2h

OpenStudy (anonymous):

@mark_o This is where i'm stuck

OpenStudy (anonymous):

f(x+h) = 4x^2+8xh+4h^2 - 2x-2h next here is f(x+h)- f(x)= 4x^2+8xh+4h^2 - 2x-2h -(4x^2 - 2x) f(x+h)- f(x)= 8xh + 4h^2 -2h correct ? now divide this to h to make f(x+h)- f(x) ----------- =? h

OpenStudy (anonymous):

= 8xh + 4h^2 -2h ______________________________ h =2h(4x + 2h^2 -2h) _____________________ h =8x + 2h^2

OpenStudy (anonymous):

@mark_o. will this be right?

OpenStudy (anonymous):

2h^2/h= 2h and 2h/h =2 therefore =?

OpenStudy (anonymous):

diff Quotient=8x +4h-2

OpenStudy (anonymous):

2). Let f(x)= x^2 - 25 and g(x) = x + 5 a. (f - g)(x)=f(x)-g(x)= (x^2 - 25) - ( x + 5) =? b) (f/g)(x)= f(x)/g(x)=(x^2 - 25)/( x + 5) =? hint try to factor the numerator then they cancel out

OpenStudy (anonymous):

a. (f - g)(x)=f(x)-g(x)= x^2 - 25 - x - 5 =? is it=x^2 -x - 30 ? ...yes or no b) (f/g)(x)= f(x)/g(x)=(x^2 - 25)/( x + 5) =(x +5)(x-5)/( x + 5) =? is it = x-5 ? ....yes or no...

OpenStudy (anonymous):

a. Yes b. yes I like the way you make it simple to understand. Thanks a lot.